### v1

# Open-Ended 3D Point Cloud Instance Segmentation

Abstract
Open-Vocab 3D Instance Segmentation methods (OV-3DIS) have recently demonstrated their ability to generalize to unseen objects. However, these methods still depend on predefined class names during testing, restricting the autonomy of agents. To mitigate this constraint, we propose a novel problem termed Open-Ended 3D Instance Segmentation (OE-3DIS), which eliminates the necessity for predefined class names during testing. Moreover, we contribute a comprehensive set of strong baselines, derived from OV-3DIS approaches and leveraging 2D Multimodal Large Language Models. To assess the performance of our OE-3DIS system, we introduce a novel Open-Ended score, evaluating both the semantic and geometric quality of predicted masks and their associated class names, alongside the standard AP score. Our approach demonstrates significant performance improvements over the baselines on the ScanNet200 and ScanNet++ datasets. Remarkably, our method surpasses the performance of Open3DIS, the current state-of-the-art method in OV-3DIS, even in the absence of ground-truth object class names.
1 Introduction
3D point cloud instance segmentation (3DIS) chen2021hierarchical ; Schult23mask3d ; ngo2023isbnet ; he2021dyco3d ; vu2022softgroup ; jiang2020pointgroup , also known as closed-vocabulary 3D instance segmentation, aims to segment all points in a point cloud into instances of classes predefined in the training set. However, this approach is less practical for scenarios where the test classes are unknown or different from the training classes. This limitation has led to the development of open-vocabulary 3D instance segmentation (OV-3DIS) nguyen2023open3dis ; takmaz2023openmask3d ; lu2023ovir ; yin2023sai3d ; yan2024maskclustering ; Huang2023Segment3D . Despite these advancements, OV-3DIS methods face several practical challenges. One major challenge is that class names must be predefined during testing, necessitating human intervention for scene understanding. This requirement significantly impedes the perception of truly autonomous agents. A potential solution is to predefined large and comprehensive vocabularies; however, this can lead to inaccuracies and significantly degrade the performance of OV-3DIS when an excessive number of vocabularies are used.
To overcome these limitations, we introduce a novel task called Open-Ended 3D Point Cloud Instance Segmentation (OE-3DIS). Unlike traditional methods, OE-3DIS does not require predefined class names during testing. Given a 3D point cloud and RGBD sequence, the system automatically generates a set of 3D masks along with their class names. To evaluate the performance of OE-3DIS methods, we propose a new metric called the OE score, in addition to the standard AP score. The OE score measures both the intersection-over-union (IoU) of the predicted instance masks with the ground-truth masks and the semantic similarity between the predicted and ground-truth class names.
To address this new and challenging task, we investigate several typical approaches: leveraging OV-3DIS methods and utilizing Multimodal Large Language Models (MLLMs). In the first approach, a list of class names is either predefined from large vocabularies or pre-extracted from image taggers. In the second approach, 2D visual tokens extracted from a pretrained 2D CLIP encoder are lifted to 3D and used as 3D visual token inputs to an LLM to predict class names. Among these methods, we find that pointwise visual token lifting to 3D for use as input to the LLM performs the best and is selected as our chosen approach. This method performs on par with state-of-the-art OV-3DIS approaches, where ground-truth classes are provided. Additionally, this approach is training-free, leveraging a pretrained 2D vision encoder (e.g., CLIP radford2021learning ) and a pretrained LLM (e.g., Vicuna chiang2023vicuna ), thereby mitigating the drawback of limited training data faced by many 3D-LLM methods.
To evaluate the performance of these methodologies, we conduct experiments on open-ended versions of two prominent 3D instance segmentation (3DIS) datasets: ScanNet200 rozenberszki2022language and ScanNet++ yeshwanthliu2023scannetpp . The results underscore the efficacy of our chosen approach over alternative baselines, demonstrating performance levels comparable to OV-3DIS methods that rely on ground-truth class names. Specifically, in ScanNet200, our approach attained an AP of 16.0, contrasting with the 22.2 AP achieved by Open3DIS, currently recognized as the state-of-the-art in OV-3DIS. However, our approach is superior in ScanNet++, where it outperforms Open3DIS by a significant margin (18.4 vs. 13.1).
In summary, the contributions of our work are:
  1. 1.
    We propose Open-Ended 3D Point Cloud Instance Segmentation (OE-3DIS), a task that segments 3D point clouds by instances and generates class names without predefined labels.
  2. 2.
    We introduce the OE score, which evaluates both the IoU of predicted instance masks with ground-truth masks and the semantic similarity of predicted and ground-truth class names.
  3. 3.
    We explore approaches for OE-3DIS, including leveraging OV-3DIS methods and Multimodal Large Language Models (MLLMs).
  4. 4.
    We present a training-free OE-3DIS method that lifts 2D visual tokens to 3D for LLM input, utilizing pretrained models like EVA-CLIP and Vicuna. This approach outperforms other baselines and achieves performance comparable to OV-3DIS methods using ground-truth class names on ScanNet200 and ScanNet++ benchmarks.
2 Related Work
3D instance segmentation (3DIS) methods such as Mask3D Schult23mask3d , ISBNet ngo2023isbnet , PointGroup jiang2020pointgroup , and others lu2023query ; robert2023spt ; robert2024scalable ; Al_Khatib_2023_spatial cluster a point cloud scene into 3D instance masks of classes predefined in the training set. These methods utilize a 3D Convolutional backbone choy2019fully ; spconv2022 ; choy2020high to extract semantic information from the 3D scene. Subsequently, they employ either Dynamic Convolution-based he2021dyco3d or Grouping-based vu2022softgroup modules to generate 3D instance masks. Recently, some approaches have adopted techniques to back-project 2D information aggregated from multiple views onto the 3D point cloud to create an ensemble of 3D point cloud features puy24scalr ; Huang2023Segment3D ; Peng2023OpenScene ; guo2023samgraph . These 2D-derived features contain rich semantic information, while those derived from 3D capture the geometrical structure of 3D objects. Combined, they supervise a 3D instance decoder to refine segmentation masks. However, these methods are closed-vocab or cannot segment new classes in testing, limiting the capability of agents to understand new 3D scenes.
Open-vocabulary 3D instance segmentation (OV-3DIS) aims at segmenting 3D objects of classes newly provided in testing. To provide 3D proposals for object recognition, OpenMask3D takmaz2023openmask3d and Lowis3D ding2023lowis3d employ 3DIS networks ngo2023isbnet ; Schult23mask3d to generate class-agnostic 3D proposals, while SAI3D yin2023sai3d , MaskClustering yan2024maskclustering , and OVIR lu2023ovir utilize 2D segmenter for producing masks for each view and consistently lift these masks to 3D. While 3DIS networks excel in capturing large geometrical structures, they often struggle to detect rare and small-shaped objects. Conversely, 2D segmenters are adept at focusing on small regions but face challenges in maintaining object consistency when lifting to the 3D point cloud. Open3DIS nguyen2023open3dis addresses these limitations by combining both 3D and 2D branches, resulting in superior results. This approach effectively captures rare and small objects while preserving the 3D geometrical structures of large objects using superpoint-level masks. While OV-3DIS is useful in some scenarios, the constraint of a predefined vocabulary set in testing requires human intervention, hindering truly autonomous agents.
3D scene understanding with Large Language Models (LLMs). Utilizing LLMs for 3D scene understanding focuses on how objects are aligned, their directions, and their locations based on textual questions within 3D environments. This approach emphasizes the spatial aspects of language understanding in three dimensions of data. Previous works 3dllm ; huang2023embodied ; scanqa ; achlioptas2020referit_3d ; chen2020scanrefer have contributed to providing 3D spatial data with language for various applications, including 3D instance and scene captioning chen2023ll3da ; chen2023end ; chen2021scan2cap ; SpaCap3D ; huang2023embodied ; 3dvista ; 3dllm , 3D visual answering questions huang2023embodied ; scanqa ; ma2022sqa3d ; parelli2023clip ; 3dvista ; 3dllm ; delitzas2023multi ; chen2023ll3da , 3D visual grounding 3dvista ; 3dllm ; huang2023embodied ; achlioptas2020referit_3d ; chen2020scanrefer ; wu2022eda and supporting embodied AI tasks like planning and reasoning 3dllm ; 3dvista ; huang2023embodied ; chen2023ll3da .
Open-ended 2D Image Understanding is an emerging task that addresses the need to recognize objects without predefined class names during training or testing. There is scant work on this task, with existing research primarily focusing on image classification nxtp ; conti2023vocabularyfree ; zhang2023recognize , object detection lin2024generateu , and instance segmentation you2023ferret ; yu2023towards ; Osprey . Standard 2D Multimodal LLMs (2D MLLMs), such as LLAVA liu2023llava , consist of a frozen vision encoder, a projector, and an LLM module. These models typically finetune either (1) the linear projector and the LLM or (2) a complex Q-Former projector. However, applying these methods for 3D scene understanding (3D-LLMs) is challenging due to the lack of sufficient 3D data and text description pairs to effectively train 3D-LLMs. In this paper, we present a novel approach to leveraging pretrained 2D MLLMs for OE-3DIS.
OmniScient Model (OSM) yu2023towards is a recently proposed pretrained 2D Multimodal LLM for Open-ended 2D Instance Segmentation, which serves as the foundation for our proposed method. OSM comprises three main modules: a visual encoder, a MaskQ-Former, and a Large Language Model (LLM). The visual encoder is a pretrained EVA-CLIP fang2023eva , a variant of the CLIP model radford2021learning , which extracts high-resolution visual features using a sliding-window scheme and incorporates global positional embeddings to preserve spatial information. The MaskQ-Former, a customized version of Q-Former instructblip designed to focus on the mask region rather than the entire image, converts visual features into fixed-length visual tokens. These tokens are then input into the Vicuna LLM chiang2023vicuna . The LLM processes these tokens to answer the question, “What is in the segmentation mask?” by outputting the object name.
In OSM, only the MaskQ-Former is trained to align visual features with the visual tokens for the LLM, while the visual encoder and LLM remain unchanged. The training datasets are large, including COCO coco , LVIS gupta2019lvis , ADE20K ade , and Cityscapes cordts2016cityscapes . This setup demonstrates a strong capability for recognizing objects without predefined class names in 2D images, inspiring us to extend this approach to 3D scene understanding.
3 Methodology
3.1 Problem Statement
Given a 3D point cloud scene $\mathbf{P}=\{\mathbf{p}_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times 6}$ consisting of $N$ points with $xyz$ coordinates and associated $rgb$ colors, along with $T$ RGB-D frames with color images $\{\mathbf{I}_{t}\}_{t=1}^{T}$ and depth ones $\{\mathbf{D}_{t}\}_{t=1}^{T}$, where $\mathbf{I}_{t}\in\mathbb{R}^{H\times W\times 3},\mathbf{D}_{t}\in\mathbb{R}_{+}^{H\times W}$, we aim to segment all $K$ object binary masks $\{\mathbf{m}_{k}\}_{k=1}^{K}$, $\mathbf{m}_{k}\in\{0,1\}^{N}$ and their associated class names $\{l_{k}\}_{k=1}^{K}$ without giving any predefined class names in testing.
The auxiliary information includes the intrinsic matrix $\mathbf{\Gamma}\in\mathbb{R}^{3\times 3}$ and the extrinsic matrix $\mathbf{[R|v]}_{t}\in\mathbb{R}^{3\times 4}$, where $H,W$ are the image’s height and width, $\mathbf{R}$ is a 3D rotation matrix, and $\mathbf{v}$ is a 3D translation vector. This composite matrix of rotation and translation converts coordinates from the global frame (of the point cloud) to the camera’s frame at view $t$.
3.2 Evaluation Metrics
To evaluate open-ended object detection or instance segmentation, where predicted class names may be similar but not exactly the same as ground-truth (GT) class names, prior work Rewatbowornwong2023ZeroGuideSeg proposed a label reassignment technique. This method uses text encoders (e.g., CLIP radford2021learning , BERT devlin2018bert , Sentence Transformer sentence-transformer to encode both the predicted and GT class names for each scene. It then matches each predicted class name to its closest GT class name based on cosine similarity. After this matching, the standard AP score is used to evaluate performance.
However, the AP score described above is based on a many-to-one matching where multiple predicted class names can be matched to the same ground-truth class name, leading to inaccuracies in fine-grained class prediction. To address this, we propose a new metric called the Open-ended score (OE score), which considers both textual semantic similarity and spatial similarity of predicted and ground-truth masks and their class names. Specifically, given $K$ predicted instances with masks, class names, and confidence scores $\{\mathbf{m}_{k}\}_{k=1}^{K}$, $\{l_{k}\}_{k=1}^{K}$, and $\{\rho_{k}\}_{k=1}^{K}$, and $J$ ground-truth instances with masks and class names as $\{\mathbf{m}_{j}\}_{j=1}^{J}$ and $\{l_{j}\}_{j=1}^{J}$, we first compute the cosine similarity $s_{k,j}$ between the text embeddings of the predicted and ground-truth class names, along with the IoU $o_{k,j}$ between predicted and ground-truth masks. Next, we construct the cost matrix $\mathbf{C}\in[0,1]^{K\times J}$ for one-to-one Hungarian matching as follows:
$\mathbf{C}(k,j)=-\sqrt[3]{o_{k,j}\cdot\rho_{k}\cdot\max(0,s_{k,j})}.$ (1)
In this way, only the best-matched 3D predictions in both text and mask are considered for evaluation. The final OE score is obtained by averaging all costs of matched pairs, or
OE $\text{OE score}=100*\frac{\sum_{(k,j)\in\text{all matched pairs}}-\mathbf{C}(k,j)}{J}.$ (2)
We average over $J$ GT masks since the number of GT masks is always smaller than the number of predicted masks.
Figure 1: Our baselines: Large-vocab (Left), Image tagging (Middle), and Maskwise (Right).
3.3 Proposed Baselines
Since OE-3DIS is very new and challenging, even more so than OV-3DIS, we focus our efforts on investigating prominent baselines. These baselines are illustrated in Fig. 1. They require a list of class-agnostic 3D mask proposals pre-extracted from Open3DIS nguyen2023open3dis with the DETIC 2D segmenter.
Large-vocab approach (Fig. 1 - Left): We start with a simple OE-3DIS baseline by using a large vocabulary of 21K common classes introduced by DETIC detic as predefined class names for OV-3DIS methods like Open3DIS nguyen2023open3dis and OpenMask3D takmaz2023openmask3d . However, this approach fails to achieve robust class prediction. This is because the fixed large vocabulary set contains multiple synonyms, resulting in uninformative class predictions after the Softmax operation.
Image Tagging approach (Fig. 1 - Middle): To reduce the number of classes, we leverage image-tagging techniques such as RAM++ zhang2023recognize to obtain only relevant class names per scene. Specifically, for each input view, a set of image tags is generated and then combined across all processed input views. The resulting unified tag set serves as the vocabulary for OV-3DIS. However, these methods often produce inconsistent class names across views, leading to redundant and similar class names.
Maskwise approach (Fig. 1 - Right): To tackle the inconsistency in class names across views, we first apply an OE-2DIS method, such as OSM yu2023towards , to each view to obtain a list of 2D masks and their predicted fixed-length visual tokens. For each 3D mask proposal, we project it onto a view and associate it with the best-matched 2D mask based on IoU to obtain its 2D fixed-length visual tokens. The 3D visual tokens for the 3D mask are then aggregated by averaging the 2D visual tokens across views, which are subsequently input into a pretrained LLM to obtain the final class names. However, this approach heavily relies on the matching process between the 3D proposal and 2D masks in each view, which is often misaligned due to imperfect segmentation and depth maps. In other words, a 3D mask can project onto multiple 2D masks, and selecting only the best-matched 2D mask might result in the loss of important information.
Figure 2: Overview of our approach. First, we generate class-agnostic 2D instance segmentation masks for all views using segmenters like DETIC detic and SAM sam , and lift these 2D masks into 3D masks using Open3DIS nguyen2023open3dis . Simultaneously, the 2D masks and their corresponding RGB images are used to extract 2D visual tokens from an MLLM like OSM yu2023towards , which are then lifted into pointwise 3D visual tokens. Finally, for each 3D proposal mask, we aggregate the pointwise 3D visual tokens to form the final tokens for input to the LLM to predict final class names.
3.4 Our Approach
To address the above limitations, we propose a method for producing pointwise 3D visual tokens, as illustrated in Fig. 2. First, we generate class-agnostic 2D instance segmentation masks for all views using class-agnostic 2D segmenters such as DETIC detic and SAM sam . Next, we lift 2D masks into 3D masks using Open3DIS nguyen2023open3dis . Simultaneously, these 2D masks, along with their corresponding RGB images, are used to extract 2D visual tokens from an MLLM like OSM yu2023towards . We then lift the resulting 2D visual tokens $\mathbf{F}^{\text{2D}}$ into 3D visual tokens to obtain pointwise 3D visual tokens $\mathbf{F}^{\text{3D}}$. Finally, for each 3D proposal mask, we query the 3D visual tokens associated with this proposal by aggregating the pointwise lifted 3D visual tokens, forming the final tokens $\mathbf{f}^{\text{3D}}$ for input to the LLM.
This approach takes into account the depth and geometric structure of 3D objects via lifting, resulting in more robust visual tokens and a unified, densely-featured point cloud that can be queried instantly at test time. Subsequently, we will focus on our pointwise visual tokens lifting and aggregation.
First, the correspondence of a 3D point $\mathbf{p}_{i}(x,y,z)\in\mathbf{P}$ with its 2D projection $(u,v)$ in view $t$ is:
$d_{i,t}\cdot\begin{bmatrix}u_{i}\\ v_{i}\\ 1\end{bmatrix}_{t}=\mathbf{\Gamma}\cdot\mathbf{[R|c]}_{t}\cdot\begin{bmatrix}x_{i}\\ y_{i}\\ z_{i}\\ 1\end{bmatrix}$ (3)
where $d_{i,t}$ is the projected 2D depth value of point $i$ to frame $t$.
For each view $t$, we extract the 2D visual tokens $\mathbf{F}^{k,\text{2D}}_{t}\in\mathbb{R}^{E\times C}$, where $E,C$ are the number of visual tokens and feature dimensions, for each 2D mask $k$ from the MaskQ-Former module of OSM yu2023towards . Furthermore, we denote $\lambda^{k}_{t,i}=\{0,1\}$ as the visibility value indicating a point $i$ is visible in mask $k$ of view $t$, $\lambda^{k}_{t}=\{0,1\}^{N}$. We set the visibility value $\lambda_{t,i}^{k}$ of any points whose pixel projections fall outside the $k$-th 2D mask’s boundaries or the disparity between projected depth $d$ and the collected depth $\mathbf{D}$ exceeds a defined depth threshold $\tau_{\text{depth}}$, or $|d_{i,t}-\mathbf{D}_{t}[\lfloor u_{i}\rfloor,\lfloor v_{i}\rfloor]|>\tau_{depth}$, to 00.
We accumulate 3D visual tokens $\mathbf{F}^{\text{3D}}\in\mathbb{R}^{N\times E\times C}$, from every 2D mask $k$ and compute for the frequency $\mathbf{r}^{\text{3D}}\in\mathbb{N}^{N}$ of every view as follows:
$\mathbf{F}^{\text{3D}}=\sum_{t,k}\lambda^{k}_{t}*\mathbf{F}^{k,\text{2D}}_{t},\quad\quad\quad\quad\quad\quad\mathbf{r}^{\text{3D}}=\sum_{t,k}\lambda^{k}_{t},$ (4)
where * is the outer product. The normalized pointwise 3D visual tokens are then obtained as follows:
$\bar{\mathbf{F}}^{\text{3D}}_{i}=\begin{cases}\mathbf{F}^{\text{3D}}_{i}/\mathbf{r}^{\text{3D}}_{i}&\text{if }\mathbf{r}^{\text{3D}}_{i}>0\\ \mathbf{0}&\text{otherwise}\end{cases}.$ (5)
Finally, for each 3D mask $\mathbf{m}_{k}$, we weighted average the visual tokens of all its points by their frequency to obtain the 3D visual tokens $\mathbf{f}^{\text{3D}}_{k}$ for that mask, which are then used as input to the LLM to predict the final class, as follows.
$\mathbf{f}^{\text{3D}}_{k}=\frac{\sum_{i\in\mathbf{m}_{k}}\bar{\mathbf{F}}^{\text{3D}}_{i}\cdot\mathbf{r}_{i}^{\text{3D}}}{\sum_{i\in\mathbf{m}_{k}}\mathbf{r}_{i}^{\text{3D}}}.$ (6)
4 Experimental Results
Datasets: We conducted experiments to assess the performance of the baselines and our proposed method using two common 3DIS datasets: ScanNet200 rozenberszki2022language and ScanNet++ yeshwanthliu2023scannetpp . The ScanNet200 dataset builds on the original ScanNet dai2017scannet by expanding its semantic categories from 20 to 200. Its instance segmentation benchmark includes 1,201 training scenes and 312 validation scenes with 198 object categories, significantly enriching the vocabulary and enhancing its capability for detailed 3D semantic and instance segmentation. The ScanNet++ dataset was recently introduced, featuring up to 1,659 semantic categories, with 360 training scenes and 50 validation scenes. Given the large number of classes, we follow the standard 3DIS evaluation protocol on ScanNet++ and evaluate only the most common 100 object categories. This dataset offers a much denser 3D point cloud scene representation, making it the most challenging dataset for 3D understanding.
[…]
Implementation details: Following Open3DIS nguyen2023open3dis , we generate class-agnostic 3D proposals from ISBNet ngo2023isbnet pretrained on ScanNet200 or by lifting 2D masks to 3D using Detic detic . For RAM++ zhang2023recognize , we employ the Swin-L model, trained on a 14-million image dataset with an image size of 384px and a tagging threshold of 0.68. For LLM, we leverage Vicuna-7B vicuna2023 , fine-tuned for open-ended 2D Instance Segmentation yu2023towards .
Table 1: Comparative results on the ScanNet200 dataset. Shaded text indicates a reference method, not a direct comparison. ‘-’ indicates results are not provided. The best results are in bold.
[…]
Mask3D Schult23mask3d 26.9 36.2 41.4 39.8 21.7 17.9 -
Fully-sup 3DIS ISBNet ngo2023isbnet 24.5 32.7 37.6 38.6 20.5 12.5 -
OpenMask3D takmaz2023openmask3d 15.4 19.9 23.1 17.1 14.1 14.9 -
OV-3DIS Open3DIS nguyen2023open3dis 23.7 29.4 32.8 27.8 21.2 21.8 -
Large-Vocab 21K DETIC classes (Ours) 8.5 11.7 13.1 9.9 7.2 8.3 52.2
Image-Tagging RAM++ zhang2023recognize (Ours) 10.7 14.3 16.0 11.6 11.0 9.3 60.5
Maskwise OSM yu2023towards (Ours) 14.4 19.8 23.9 18.9 13.5 10.2 65.0
Pointwise Ours 16.0 22.0 24.7 20.0 14.3 13.2 65.2
Table 2: Comparative results on the ScanNet++ dataset. Shaded text indicates a reference method, not a direct comparison. ‘-’ indicates results are not provided. The best results are in bold.
Setting Method AP AP50 AP25 AR RC50 RC25 OE
PointGroup jiang2020pointgroup 8.9 14.6 38.9 - - - -
Fully-sup 3DIS SoftGroup vu2022softgroup 16.7 29.7 21.0 - - - -
OpenMask3D takmaz2023openmask3d 2.0 2.7 3.4 4.6 8.3 12.4 -
OV-3DIS Open3DIS (known classes) nguyen2023open3dis 13.1 20.8 24.6 22.1 33.9 39.1 -
Large-Vocab 21K DETIC classes (Ours) 7.3 11.9 15.2 13.3 20.3 23.6 42.0
Image-Tagging RAM++ zhang2023recognize (Ours) 9.1 15.5 19.1 16.0 24.8 28.7 48.5
Maskwise OSM yu2023towards (Ours) 16.3 24.8 29.0 22.2 32.0 36.0 51.2
Pointwise Ours 18.4 29.4 33.6 23.3 35.2 39.3 53.9
4.1 Comparison with Baselines
We compare our approach to the proposed baselines using the OE-3DIS setting on the ScanNet200 (Tab. 1) and ScanNet++ (Tab. 2) datasets. For reference, we also present results from OV-3DIS methods, including OpenMask3D takmaz2023openmask3d and Open3DIS nguyen2023open3dis , as well as fully-supervised methods like ISBNet ngo2023isbnet and Mask3D Schult23mask3d for ScanNet200; and PointGroup jiang2020pointgroup and SoftGroup vu2022softgroup for ScanNet++.
For Scannet200: We obtain class-agnostic 3D proposals from two sources: 3D masks from a 3DIS network such as ISBNet ngo2023isbnet , and 2D Lift 3D masks from Open3DIS nguyen2023open3dis . Our proposed approach outperforms other baselines in both AP and OE scores. Furthermore, the performance progression of the baselines, in the specified order, clearly justifies the motivation behind each baseline compared to its predecessor, as discussed in Sec. 3.3. Interestingly, our approach also surpasses OpenMask3D takmaz2023openmask3d (16.0 vs. 15.4 in AP), even though OpenMask3D utilizes provided class names. This indicates that, in some cases, we can achieve OV-3DIS without relying on provided class names.
For ScanNet++: Due to the extensive scale and vast array of classes in ScanNet++, the performance of 3D mask results from ISBNet is inadequate. Consequently, we solely rely on the utilization of 2D Lift 3D masks from Open3DIS nguyen2023open3dis . We notice a consistent trend akin to the results observed in ScanNet200. Particularly noteworthy is the significant outperformance of our approach compared to OV-3DIS methods or even fully-supervised 3DIS, as evidenced by higher AP scores. This underscores the promising application of OE-3DIS in navigating complex 3D scene structures.
Qualitative comparison: In Fig. 3, both our proposed baselines and approach effectively assign class names to 3D proposals in OE-3DIS settings. The first column of ScanNet200 qualitative results shows our architectures accurately predicting all 3D proposals. The second column highlights differences between our approach and the baselines. Our method precisely identifies the ‘painting’ object as “wall painting”, while other baselines predict less accurate object labels. In the final column, our 3D mask proposals offer more detail than ground truth, enabling predictions of classes not in ScanNet++’s vocabulary, such as “photocopier”.
Figure 3: Qualitative examples are provided for ScanNet200 dai2017scannet (first two columns) and ScanNet++ yeshwanthliu2023scannetpp (last two columns). Both our baselines and approach yield notably good results, particularly in terms of accurately identifying class names even though they do not match exactly the GT classes.
Table 3: Ablation on point aggregation techniques
[…]
Max 6.3 8.4 10.2 8.5 4.7 5.6 56.3
Random 13.2 18.9 21.6 17.7 12.3 10.0 60.4
Mean 14.5 20.1 22.6 18.2 13.1 11.2 63.1
Weighted Average 16.0 22.0 24.7 20.0 14.3 13.2 65.2
Table 4: Study on different types 3D proposals.
[…]
3D masks DETIC 11.7 16.0 18.3 16.3 10.3 8.1 56.4
2D Lift 3D masks DETIC 11.7 18.6 23.5 11.3 12.1 11.8 54.3
3D + 2D Lift 3D masks DETIC 16.0 22.0 24.7 20.0 14.3 13.2 65.2
3D + 2D Lift 3D masks SAM 15.4 20.5 22.9 19.2 14.8 11.7 64.8
4.2 Ablation Study
To study many design choices of our pointwise approach, we intensively carry out ablation study on the ScanNet200 dai2017scannet dataset.
Study on different point feature aggregation techniques (Eq. 6) is shown in Tab. 3. We evaluate four techniques for combining point features from a given 3D mask: max, random (randomly selecting one point), mean, and weighted average (our proposed operation). The weighted average technique achieved the highest performance with an AP score of 16.0, outperforming the other methods. Consequently, we utilized the weighted average technique in all our experiments.
Study on different types 3D proposals. As reported in Tab. 4, combining 3D proposals from both 3D masks and 2D Lift 3D masks yielded the most favorable results. This outcome validates our choice of 3D proposals for the ScanNet200 dataset. Additionally, the 2D Lift 3D proposals obtained from DETIC demonstrated slightly superior performance compared to those from SAM. This observation diverges from the methodology adopted in Open3DIS nguyen2023open3dis , where provided class names are utilized to employ Grounded-SAM ren2024grounded instead of SAM sam , as in our OE-3DIS scenario.
Study on different text encoders for evaluation metrics is described in Tab. 5. We assess three text decoders: BERT devlin2018bert , CLIP radford2021learning , and Sentence Transformer sentence-transformer . Among them, the Sentence Transformer embeddings attain the highest AP scores, whereas the CLIP text embeddings achieve the highest OE score. The OE score indicates that the text embeddings from CLIP exhibit a higher degree of similarity to each other, contrasting with those from the Sentence Transformer, which tend to describe complex sentences rather than solely focusing on a single class name. Hence, we lean towards selecting the Sentence Transformer as our recommended text encoder.
Study on different text prompts used to query the LLM: We experiment with various input text prompts to query the LLM for class names. Tab. 6 demonstrates that altering the prompt subtly affects the model’s accuracy. We select the prompt “What is in the segmentation mask? Assistant:” from the variants for our approach.
Table 5: Study on different text encoders used in evaluation metrics
[…]
BERT devlin2018bert 13.5 19.0 21.2 17.6 11.6 11.0 70.6
CLIP radford2021learning 15.9 22.0 24.5 20.3 13.7 13.3 71.6
Sentence Transformer sentence-transformer 16.0 22.0 24.7 20.0 14.3 13.2 65.2
Table 6: Study on different text prompts used to query the LLM.
Text Prompt AP OE
“What is in the segmentation mask? Assistant:” 16.0 65.2
“Can you describe what is in the segmentation mask region? Assistant:” 15.3 64.3
“What can you see in the segmentation mask region? Assistant:” 15.9 65.0
“Could you use a few words to describe what is in the segmentation mask region? Assistant:” 15.2 64.0
“What is this segmentation mask? Assistant:” 15.8 65.0
5 Discussion and Conclusion
Limitations: Though our method enhances open-ended 3D scene understanding, it has notable drawbacks. Firstly, it heavily relies on 2D visual tokens from the pretrained OSM, which are trained on instance segmentation datasets with limited classes. This hampers the model’s ability to recognize a wide range of classes in an open-world environment. Secondly, the performance of OE-3DIS is contingent upon class-agnostic 3D instance segmentation (3DIS) methods, which in turn rely on the quality of 3D representation and 2D-3D mapping, such as camera calibrations and depth images.
Conclusion: We have introduced Open-Ended 3D Point Cloud Instance Segmentation (OE-3DIS), which generates 3D masks and object class names without predefined labels during testing. We have explored baselines using OV-3DIS methods and MLLMs, and introduced a pointwise training-free approach leveraging OSM. We have also proposed the OE score, tailored for OE-3DIS. Experiments on ScanNet200 and ScanNet++ show our approach’s superior performance, notably outperforming Open3DIS (SOTA on OV-3DIS) on ScanNet++ without ground-truth class names.

### v2

@@
 Abstract
-Open-Vocab 3D Instance Segmentation methods (OV-3DIS) have recently demonstrated their ability to generalize to unseen objects.
-However, these methods still depend on predefined class names during testing, restricting the autonomy of agents.
+Open-vocabulary 3D Instance Segmentation methods (OV-3DIS) have recently demonstrated their generalization ability to unseen objects.
+However, these methods still depend on predefined class names during inference, restricting agents’ autonomy.
 To mitigate this constraint, we propose a novel problem termed Open-Ended 3D Instance Segmentation (OE-3DIS), which eliminates the necessity for predefined class names during testing.
-Moreover, we contribute a comprehensive set of strong baselines, derived from OV-3DIS approaches and leveraging 2D Multimodal Large Language Models.
-To assess the performance of our OE-3DIS system, we introduce a novel Open-Ended score, evaluating both the semantic and geometric quality of predicted masks and their associated class names, alongside the standard AP score.
-Our approach demonstrates significant performance improvements over the baselines on the ScanNet200 and ScanNet++ datasets.
-Remarkably, our method surpasses the performance of Open3DIS, the current state-of-the-art method in OV-3DIS, even in the absence of ground-truth object class names.
+We present a comprehensive set of strong baselines inspired by OV-3DIS methodologies, utilizing 2D Multimodal Large Language Models.
+In addition, we introduce a novel token aggregation strategy that effectively fuses information from multiview images.
+To evaluate the performance of our OE-3DIS system, we benchmark both the proposed baselines and our method on two widely used indoor datasets: ScanNet200 and ScanNet++.
+Our approach achieves substantial performance gains over the baselines on both datasets.
+Notably, even without access to ground-truth object class names during inference, our method outperforms Open3DIS, the current state-of-the-art in OV-3DIS.
+Source code available at: https://github.com/PhucNDA/OE-3DIS.
+Figure 1: During testing, Open-Vocabulary 3D Instance Segmentation (OV-3DIS) utilizes CLIP to generate confidence scores for provided text prompts.
+However, its reliance on a predefined vocabulary during inference necessitates human intervention, restricting autonomy in intelligent agents.
+In contrast, our proposed Open-Ended 3D Instance Segmentation (OE-3DIS) leverages Large Language Models to enable open interaction with humans, recognizing fine-grained objects without being constrained by predefined class labels.
+**footnotetext: These authors contributed equally to this work
 1 Introduction
-3D point cloud instance segmentation (3DIS) chen2021hierarchical ; Schult23mask3d ; ngo2023isbnet ; he2021dyco3d ; vu2022softgroup ; jiang2020pointgroup , also known as closed-vocabulary 3D instance segmentation, aims to segment all points in a point cloud into instances of classes predefined in the training set.
+3D point cloud instance segmentation (3DIS) [5, 50, 37, 25, 52, 30], also known as closed-vocabulary 3D instance segmentation, aims to segment all points in a point cloud into instances of classes predefined in the training set.
 However, this approach is less practical for scenarios where the test classes are unknown or different from the training classes.
-This limitation has led to the development of open-vocabulary 3D instance segmentation (OV-3DIS) nguyen2023open3dis ; takmaz2023openmask3d ; lu2023ovir ; yin2023sai3d ; yan2024maskclustering ; Huang2023Segment3D .
+This limitation has led to the development of open-vocabulary 3D instance segmentation (OV-3DIS) [39, 51, 35, 57, 55, 28].
 Despite these advancements, OV-3DIS methods face several practical challenges.
-One major challenge is that class names must be predefined during testing, necessitating human intervention for scene understanding.
+One major challenge is that class names must be predefined during inference, necessitating human intervention for scene understanding.
 This requirement significantly impedes the perception of truly autonomous agents.
 A potential solution is to predefined large and comprehensive vocabularies; however, this can lead to inaccuracies and significantly degrade the performance of OV-3DIS when an excessive number of vocabularies are used.
-To overcome these limitations, we introduce a novel task called Open-Ended 3D Point Cloud Instance Segmentation (OE-3DIS).
+To overcome these limitations, we introduce a novel task called Open-Ended 3D Point Cloud Instance Segmentation (OE-3DIS) see Fig.
+1.
 Unlike traditional methods, OE-3DIS does not require predefined class names during testing.
 Given a 3D point cloud and RGBD sequence, the system automatically generates a set of 3D masks along with their class names.
-To evaluate the performance of OE-3DIS methods, we propose a new metric called the OE score, in addition to the standard AP score.
-The OE score measures both the intersection-over-union (IoU) of the predicted instance masks with the ground-truth masks and the semantic similarity between the predicted and ground-truth class names.
-To address this new and challenging task, we investigate several typical approaches: leveraging OV-3DIS methods and utilizing Multimodal Large Language Models (MLLMs).
-In the first approach, a list of class names is either predefined from large vocabularies or pre-extracted from image taggers.
-In the second approach, 2D visual tokens extracted from a pretrained 2D CLIP encoder are lifted to 3D and used as 3D visual token inputs to an LLM to predict class names.
-Among these methods, we find that pointwise visual token lifting to 3D for use as input to the LLM performs the best and is selected as our chosen approach.
-This method performs on par with state-of-the-art OV-3DIS approaches, where ground-truth classes are provided.
-Additionally, this approach is training-free, leveraging a pretrained 2D vision encoder (e.g., CLIP radford2021learning ) and a pretrained LLM (e.g., Vicuna chiang2023vicuna ), thereby mitigating the drawback of limited training data faced by many 3D-LLM methods.
-To evaluate the performance of these methodologies, we conduct experiments on open-ended versions of two prominent 3D instance segmentation (3DIS) datasets: ScanNet200 rozenberszki2022language and ScanNet++ yeshwanthliu2023scannetpp .
+To evaluate the performance of OE-3DIS methods, we follow [46] to leverage an adaptive protocol of standard AP score calculation.
+To tackle this new and challenging task, we introduce a novel method along with our proposed three established baselines that leverage OV-3DIS techniques and Multimodal Large Language Models (MLLMs).
+In the first two baselines, class names are either predefined from extensive vocabularies or extracted using image taggers.
+The third baseline employs 2D visual tokens, obtained from a pretrained 2D CLIP encoder, which are then processed through a Visual Token Aggregation mechanism before being fed into an LLM to predict class names.
+Finally, our proposed method takes this further by consistently aggregating 2D visual tokens from multiview images, transforming them into dense 3D point cloud tokens that enable real-time querying.
+Notably, our pointwise visual token lifting approach, which feeds these tokens into an LLM, achieves the best performance across multiple benchmarks, matching state-of-the-art OV-3DIS methods that rely on ground-truth class labels.
+Furthermore, both our proposed method and designed baselines are entirely training-free, utilizing only pretrained 2D vision encoders (e.g., CLIP [43]) and pretrained LLMs (e.g., Vicuna [10]).
+The training-free approach effectively mitigates the prevalent issue of insufficient training data that hampers many existing 3D-LLM techniques.
+To evaluate the performance of these methods, we conduct experiments on open-ended versions of two prominent 3D instance segmentation (3DIS) datasets: ScanNet200 [49] and ScanNet++ [56].
 The results underscore the efficacy of our chosen approach over alternative baselines, demonstrating performance levels comparable to OV-3DIS methods that rely on ground-truth class names.
@@
 However, our approach is superior in ScanNet++, where it outperforms Open3DIS by a significant margin (18.4 vs.
-13.1).
+13.1 in AP).
 In summary, the contributions of our work are:
@@
   2. 2.
-    We introduce the OE score, which evaluates both the IoU of predicted instance masks with ground-truth masks and the semantic similarity of predicted and ground-truth class names.
+    We establish solid baselines for OE-3DIS, including leveraging OV-3DIS methods and Multimodal Large Language Models (MLLMs).
   3. 3.
-    We explore approaches for OE-3DIS, including leveraging OV-3DIS methods and Multimodal Large Language Models (MLLMs).
-  4. 4.
-    We present a training-free OE-3DIS method that lifts 2D visual tokens to 3D for LLM input, utilizing pretrained models like EVA-CLIP and Vicuna.
-This approach outperforms other baselines and achieves performance comparable to OV-3DIS methods using ground-truth class names on ScanNet200 and ScanNet++ benchmarks.
+    We present a training-free OE-3DIS method that lifts 2D visual tokens to 3D and utilizes pretrained Multimodal LLMs to output the final object classes.
 2 Related Work
-3D instance segmentation (3DIS) methods such as Mask3D Schult23mask3d , ISBNet ngo2023isbnet , PointGroup jiang2020pointgroup , and others lu2023query ; robert2023spt ; robert2024scalable ; Al_Khatib_2023_spatial cluster a point cloud scene into 3D instance masks of classes predefined in the training set.
-These methods utilize a 3D Convolutional backbone choy2019fully ; spconv2022 ; choy2020high to extract semantic information from the 3D scene.
-Subsequently, they employ either Dynamic Convolution-based he2021dyco3d or Grouping-based vu2022softgroup modules to generate 3D instance masks.
-Recently, some approaches have adopted techniques to back-project 2D information aggregated from multiple views onto the 3D point cloud to create an ensemble of 3D point cloud features puy24scalr ; Huang2023Segment3D ; Peng2023OpenScene ; guo2023samgraph .
+3D instance segmentation (3DIS) methods such as Mask3D [50], ISBNet [37], PointGroup [30], and others [34, 47, 48, 2] cluster a point cloud scene into 3D instance masks of classes predefined in the training set.
+These methods utilize a 3D Convolutional backbone [11, 15, 12] to extract semantic information from the 3D scene.
+Subsequently, they employ either Dynamic Convolution-based [25] or Grouping-based [52] modules to generate 3D instance masks.
+Recently, some approaches have adopted techniques to back-project 2D information aggregated from multiple views onto the 3D point cloud to create an ensemble of 3D point cloud features [42, 28, 41, 23].
 These 2D-derived features contain rich semantic information, while those derived from 3D capture the geometrical structure of 3D objects.
 Combined, they supervise a 3D instance decoder to refine segmentation masks.
-However, these methods are closed-vocab or cannot segment new classes in testing, limiting the capability of agents to understand new 3D scenes.
+However, these methods are closed-vocab or cannot segment new classes in testing, limiting the capability to understand new 3D scenes.
 Open-vocabulary 3D instance segmentation (OV-3DIS) aims at segmenting 3D objects of classes newly provided in testing.
-To provide 3D proposals for object recognition, OpenMask3D takmaz2023openmask3d and Lowis3D ding2023lowis3d employ 3DIS networks ngo2023isbnet ; Schult23mask3d to generate class-agnostic 3D proposals, while SAI3D yin2023sai3d , MaskClustering yan2024maskclustering , and OVIR lu2023ovir utilize 2D segmenter for producing masks for each view and consistently lift these masks to 3D.
+To provide 3D proposals for object recognition, OpenMask3D [51] and Lowis3D [21] employ 3DIS networks [37, 50] to generate class-agnostic 3D proposals, while SAI3D [57], MaskClustering [55], OVIR [35] and Any3DIS [38] utilize 2D segmenter for producing masks for each view and lift these masks to 3D.
 While 3DIS networks excel in capturing large geometrical structures, they often struggle to detect rare and small-shaped objects.
 Conversely, 2D segmenters are adept at focusing on small regions but face challenges in maintaining object consistency when lifting to the 3D point cloud.
-Open3DIS nguyen2023open3dis addresses these limitations by combining both 3D and 2D branches, resulting in superior results.
+Open3DIS [39] addresses these limitations by combining both 3D and 2D branches, resulting in superior results.
 This approach effectively captures rare and small objects while preserving the 3D geometrical structures of large objects using superpoint-level masks.
-While OV-3DIS is useful in some scenarios, the constraint of a predefined vocabulary set in testing requires human intervention, hindering truly autonomous agents.
+While OV-3DIS is useful in some scenarios, the constraint of a predefined vocabulary set in inference requires human intervention, hindering very autonomous agents.
 3D scene understanding with Large Language Models (LLMs).
@@
 This approach emphasizes the spatial aspects of language understanding in three dimensions of data.
-Previous works 3dllm ; huang2023embodied ; scanqa ; achlioptas2020referit_3d ; chen2020scanrefer have contributed to providing 3D spatial data with language for various applications, including 3D instance and scene captioning chen2023ll3da ; chen2023end ; chen2021scan2cap ; SpaCap3D ; huang2023embodied ; 3dvista ; 3dllm , 3D visual answering questions huang2023embodied ; scanqa ; ma2022sqa3d ; parelli2023clip ; 3dvista ; 3dllm ; delitzas2023multi ; chen2023ll3da , 3D visual grounding 3dvista ; 3dllm ; huang2023embodied ; achlioptas2020referit_3d ; chen2020scanrefer ; wu2022eda and supporting embodied AI tasks like planning and reasoning 3dllm ; 3dvista ; huang2023embodied ; chen2023ll3da .
+Previous works [26, 27, 3, 1, 4] have contributed to providing 3D spatial data with language for various applications, including 3D instance and scene captioning [6, 7, 8, 53, 27, 64, 26], 3D visual answering questions [27, 3, 36, 40, 64, 26, 19, 6], 3D visual grounding [64, 26, 27, 1, 4, 54] and supporting embodied AI tasks like planning and reasoning [26, 64, 27, 6].
 Open-ended 2D Image Understanding is an emerging task that addresses the need to recognize objects without predefined class names during training or testing.
-There is scant work on this task, with existing research primarily focusing on image classification nxtp ; conti2023vocabularyfree ; zhang2023recognize , object detection lin2024generateu , and instance segmentation you2023ferret ; yu2023towards ; Osprey .
-Standard 2D Multimodal LLMs (2D MLLMs), such as LLAVA liu2023llava , consist of a frozen vision encoder, a projector, and an LLM module.
+There is scant work on this task, with existing research primarily focusing on image classification [61, 14, 29], object detection [13], and instance segmentation [58, 59, 60].
+Standard 2D Multimodal LLMs (2D MLLMs), such as LLAVA [33], consist of a frozen vision encoder, a projector, and an LLM module.
 These models typically finetune either (1) the linear projector and the LLM or (2) a complex Q-Former projector.
@@
 In this paper, we present a novel approach to leveraging pretrained 2D MLLMs for OE-3DIS.
-OmniScient Model (OSM) yu2023towards is a recently proposed pretrained 2D Multimodal LLM for Open-ended 2D Instance Segmentation, which serves as the foundation for our proposed method.
+OmniScient Model (OSM) [59] is a recently proposed pretrained 2D Multimodal LLM for Open-ended 2D Instance Segmentation, which serves as the foundation for our proposed method.
 OSM comprises three main modules: a visual encoder, a MaskQ-Former, and a Large Language Model (LLM).
-The visual encoder is a pretrained EVA-CLIP fang2023eva , a variant of the CLIP model radford2021learning , which extracts high-resolution visual features using a sliding-window scheme and incorporates global positional embeddings to preserve spatial information.
-The MaskQ-Former, a customized version of Q-Former instructblip designed to focus on the mask region rather than the entire image, converts visual features into fixed-length visual tokens.
-These tokens are then input into the Vicuna LLM chiang2023vicuna .
+The visual encoder is a pretrained EVA-CLIP [22], a variant of the CLIP model [43], which extracts high-resolution visual features using a sliding-window scheme and incorporates global positional embeddings to preserve spatial information.
+The MaskQ-Former, a customized version of Q-Former [18] designed to focus on the mask region rather than the entire image, converts visual features into fixed-length visual tokens.
+These tokens are then input into the Vicuna LLM [10].
 The LLM processes these tokens to answer the question, “What is in the segmentation mask?” by outputting the object name.
 In OSM, only the MaskQ-Former is trained to align visual features with the visual tokens for the LLM, while the visual encoder and LLM remain unchanged.
-The training datasets are large, including COCO coco , LVIS gupta2019lvis , ADE20K ade , and Cityscapes cordts2016cityscapes .
+The training datasets are large, including COCO [32], LVIS [24], ADE20K [62], and Cityscapes [16].
 This setup demonstrates a strong capability for recognizing objects without predefined class names in 2D images, inspiring us to extend this approach to 3D scene understanding.
+Figure 2: Our baselines: Large-vocab (Left), Image tagging (Middle), and Maskwise (Right).
 3 Methodology
 3.1 Problem Statement
-Given a 3D point cloud scene $\mathbf{P}=\{\mathbf{p}_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times 6}$ consisting of $N$ points with $xyz$ coordinates and associated $rgb$ colors, along with $T$ RGB-D frames with color images $\{\mathbf{I}_{t}\}_{t=1}^{T}$ and depth ones $\{\mathbf{D}_{t}\}_{t=1}^{T}$, where $\mathbf{I}_{t}\in\mathbb{R}^{H\times W\times 3},\mathbf{D}_{t}\in\mathbb{R}_{+}^{H\times W}$, we aim to segment all $K$ object binary masks $\{\mathbf{m}_{k}\}_{k=1}^{K}$, $\mathbf{m}_{k}\in\{0,1\}^{N}$ and their associated class names $\{l_{k}\}_{k=1}^{K}$ without giving any predefined class names in testing.
-The auxiliary information includes the intrinsic matrix $\mathbf{\Gamma}\in\mathbb{R}^{3\times 3}$ and the extrinsic matrix $\mathbf{[R|v]}_{t}\in\mathbb{R}^{3\times 4}$, where $H,W$ are the image’s height and width, $\mathbf{R}$ is a 3D rotation matrix, and $\mathbf{v}$ is a 3D translation vector.
-This composite matrix of rotation and translation converts coordinates from the global frame (of the point cloud) to the camera’s frame at view $t$.
+Given a 3D point cloud scene $\mathbf{P}=\{\mathbf{p}_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times 6}$ consisting of $N$ points with xyzxyz coordinates and associated rgbrgb colors, along with $T$ RGB-D frames with color images $\{\mathbf{I}_{t}\}_{t=1}^{T}$ and depth ones $\{\mathbf{D}_{t}\}_{t=1}^{T}$, where $\mathbf{I}_{t}\in\mathbb{R}^{H\times W\times 3},\mathbf{D}_{t}\in\mathbb{R}_{+}^{H\times W}$, we aim to segment all $K$ object binary masks $\{\mathbf{m}_{k}\}_{k=1}^{K}$, $\mathbf{m}_{k}\in\{0,1\}^{N}$ and their associated class names $\{l_{k}\}_{k=1}^{K}$ without giving any predefined class names in inference.
+Of course, we do need GT class names during evaluation to assess our method’s performance.
+The information includes the intrinsic $\mathbf{\Gamma}\in\mathbb{R}^{3\times 3}$ and the extrinsic $\mathbf{[R|v]}_{t}\in\mathbb{R}^{3\times 4}$.
+In this context, $H$ and $W$ represent the height and width of the image, respectively.
+The matrix $\mathbf{R}$ is a 3D rotation matrix, while $\mathbf{v}$ is a 3D translation vector.
+This composite matrix, which combines rotation and translation, converts coordinates from the global frame of the point cloud to the camera’s frame at view $t$.
 3.2 Evaluation Metrics
-To evaluate open-ended object detection or instance segmentation, where predicted class names may be similar but not exactly the same as ground-truth (GT) class names, prior work Rewatbowornwong2023ZeroGuideSeg proposed a label reassignment technique.
-This method uses text encoders (e.g., CLIP radford2021learning , BERT devlin2018bert , Sentence Transformer sentence-transformer to encode both the predicted and GT class names for each scene.
+To evaluate open-ended object detection or instance segmentation, where predicted class names may be similar but not exactly the same as ground-truth (GT) class names, prior work [46] proposed a label reassignment technique.
+This method uses text encoders (e.g., CLIP [43], BERT [20], Sentence Transformer [44] to encode both the predicted and GT class names for each scene.
 It then matches each predicted class name to its closest GT class name based on cosine similarity.
 After this matching, the standard AP score is used to evaluate performance.
-However, the AP score described above is based on a many-to-one matching where multiple predicted class names can be matched to the same ground-truth class name, leading to inaccuracies in fine-grained class prediction.
-To address this, we propose a new metric called the Open-ended score (OE score), which considers both textual semantic similarity and spatial similarity of predicted and ground-truth masks and their class names.
-Specifically, given $K$ predicted instances with masks, class names, and confidence scores $\{\mathbf{m}_{k}\}_{k=1}^{K}$, $\{l_{k}\}_{k=1}^{K}$, and $\{\rho_{k}\}_{k=1}^{K}$, and $J$ ground-truth instances with masks and class names as $\{\mathbf{m}_{j}\}_{j=1}^{J}$ and $\{l_{j}\}_{j=1}^{J}$, we first compute the cosine similarity $s_{k,j}$ between the text embeddings of the predicted and ground-truth class names, along with the IoU $o_{k,j}$ between predicted and ground-truth masks.
-Next, we construct the cost matrix $\mathbf{C}\in[0,1]^{K\times J}$ for one-to-one Hungarian matching as follows:
-$\mathbf{C}(k,j)=-\sqrt[3]{o_{k,j}\cdot\rho_{k}\cdot\max(0,s_{k,j})}.$ (1)
-In this way, only the best-matched 3D predictions in both text and mask are considered for evaluation. The final OE score is obtained by averaging all costs of matched pairs, or
-OE $\text{OE score}=100*\frac{\sum_{(k,j)\in\text{all matched pairs}}-\mathbf{C}(k,j)}{J}.$ (2)
-We average over $J$ GT masks since the number of GT masks is always smaller than the number of predicted masks.
-Figure 1: Our baselines: Large-vocab (Left), Image tagging (Middle), and Maskwise (Right).
 3.3 Proposed Baselines
-Since OE-3DIS is very new and challenging, even more so than OV-3DIS, we focus our efforts on investigating prominent baselines.
+Since OE-3DIS is very new and challenging, we focus our efforts on investigating prominent baselines.
 These baselines are illustrated in Fig.
-1.
-They require a list of class-agnostic 3D mask proposals pre-extracted from Open3DIS nguyen2023open3dis with the DETIC 2D segmenter.
+2.
+They require a list of class-agnostic 3D mask proposals pre-extracted from Open3DIS [39] with the DETIC 2D segmenter.
 Large-vocab approach (Fig.
-1 - Left): We start with a simple OE-3DIS baseline by using a large vocabulary of 21K common classes introduced by DETIC detic as predefined class names for OV-3DIS methods like Open3DIS nguyen2023open3dis and OpenMask3D takmaz2023openmask3d .
+2 - Left): We start with a simple OE-3DIS baseline by using a large vocabulary of 21K common classes introduced by DETIC [63] as predefined class names for OV-3DIS methods like Open3DIS [39] and OpenMask3D [51].
 However, this approach fails to achieve robust class prediction.
@@
 Image Tagging approach (Fig.
-1 - Middle): To reduce the number of classes, we leverage image-tagging techniques such as RAM++ zhang2023recognize to obtain only relevant class names per scene.
+2 - Middle): To reduce the number of classes, we leverage image-tagging techniques such as RAM++ [29] to obtain only relevant class names per scene.
 Specifically, for each input view, a set of image tags is generated and then combined across all processed input views.
@@
 Maskwise approach (Fig.
-1 - Right): To tackle the inconsistency in class names across views, we first apply an OE-2DIS method, such as OSM yu2023towards , to each view to obtain a list of 2D masks and their predicted fixed-length visual tokens.
+2 - Right): To tackle the inconsistency in class names across views, we first apply an OE-2DIS method, such as OSM [59], to each view to obtain a list of 2D masks and their predicted fixed-length visual tokens.
 For each 3D mask proposal, we project it onto a view and associate it with the best-matched 2D mask based on IoU to obtain its 2D fixed-length visual tokens.
 The 3D visual tokens for the 3D mask are then aggregated by averaging the 2D visual tokens across views, which are subsequently input into a pretrained LLM to obtain the final class names.
-However, this approach heavily relies on the matching process between the 3D proposal and 2D masks in each view, which is often misaligned due to imperfect segmentation and depth maps.
-In other words, a 3D mask can project onto multiple 2D masks, and selecting only the best-matched 2D mask might result in the loss of important information.
+However, this approach relies on matching 3D proposals with 2D masks, which is often misaligned due to segmentation and depth map imperfections.
+A 3D mask can project onto multiple 2D masks, and selecting only the best match may discard valuable information.
 Figure 3: Overview of our approach.
-First, we generate class-agnostic 2D instance segmentation masks for all views using segmenters like DETIC detic and SAM sam , and lift these 2D masks into 3D masks using Open3DIS nguyen2023open3dis .
-Simultaneously, the 2D masks and their corresponding RGB images are used to extract 2D visual tokens from an MLLM like OSM yu2023towards , which are then lifted into pointwise 3D visual tokens.
+First, we generate class-agnostic 2D instance segmentation masks for all views using segmenters like DETIC [63] and SAM [31], and lift these 2D masks into 3D masks using Open3DIS [39].
+Simultaneously, the 2D masks and their corresponding RGB images are used to extract 2D visual tokens from an MLLM like OSM [59], which are then lifted into pointwise 3D visual tokens.
 Finally, for each 3D proposal mask, we aggregate the pointwise 3D visual tokens to form the final tokens for input to the LLM to predict final class names.
@@
 To address the above limitations, we propose a method for producing pointwise 3D visual tokens, as illustrated in Fig.
-2.
-First, we generate class-agnostic 2D instance segmentation masks for all views using class-agnostic 2D segmenters such as DETIC detic and SAM sam .
-Next, we lift 2D masks into 3D masks using Open3DIS nguyen2023open3dis .
-Simultaneously, these 2D masks, along with their corresponding RGB images, are used to extract 2D visual tokens from an MLLM like OSM yu2023towards .
+3.
+First, we generate class-agnostic 2D instance segmentation masks for all views using class-agnostic 2D segmenters such as DETIC [63] and SAM [31].
+Next, we lift 2D masks into 3D masks using Open3DIS [39].
+Simultaneously, these 2D masks, along with their corresponding RGB images, are used to extract 2D visual tokens from an MLLM like OSM [59].
 We then lift the resulting 2D visual tokens $\mathbf{F}^{\text{2D}}$ into 3D visual tokens to obtain pointwise 3D visual tokens $\mathbf{F}^{\text{3D}}$.
@@
 Subsequently, we will focus on our pointwise visual tokens lifting and aggregation.
-First, the correspondence of a 3D point $\mathbf{p}_{i}(x,y,z)\in\mathbf{P}$ with its 2D projection $(u,v)$ in view $t$ is:
+Concretely, first, the correspondence of a 3D point $\mathbf{p}_{i}(x,y,z)\in\mathbf{P}$ with its 2D projection $(u,v)$ in view $t$ is:
 $d_{i,t}\cdot\begin{bmatrix}u_{i}\\ v_{i}\\ 1\end{bmatrix}_{t}=\mathbf{\Gamma}\cdot\mathbf{[R|c]}_{t}\cdot\begin{bmatrix}x_{i}\\ y_{i}\\ z_{i}\\ 1\end{bmatrix}$ (1)
-where $d_{i,t}$ is the projected 2D depth value of point $i$ to frame $t$.
-For each view $t$, we extract the 2D visual tokens $\mathbf{F}^{k,\text{2D}}_{t}\in\mathbb{R}^{E\times C}$, where $E,C$ are the number of visual tokens and feature dimensions, for each 2D mask $k$ from the MaskQ-Former module of OSM yu2023towards .
+where $d_{i,t}$ is the projected depth of point $i$ to frame $t$.
+Next, for each view $t$, we extract the 2D visual tokens $\mathbf{F}^{k,\text{2D}}_{t}\in\mathbb{R}^{E\times C}$, where $E,C$ are the number of visual tokens and feature dimensions, for each 2D mask $k$ from the MaskQ-Former module of OSM [59].
 Furthermore, we denote $\lambda^{k}_{t,i}=\{0,1\}$ as the visibility value indicating a point $i$ is visible in mask $k$ of view $t$, $\lambda^{k}_{t}=\{0,1\}^{N}$.
-We set the visibility value $\lambda_{t,i}^{k}$ of any points whose pixel projections fall outside the $k$-th 2D mask’s boundaries or the disparity between projected depth $d$ and the collected depth $\mathbf{D}$ exceeds a defined depth threshold $\tau_{\text{depth}}$, or $|d_{i,t}-\mathbf{D}_{t}[\lfloor u_{i}\rfloor,\lfloor v_{i}\rfloor]|>\tau_{depth}$, to 00.
-We accumulate 3D visual tokens $\mathbf{F}^{\text{3D}}\in\mathbb{R}^{N\times E\times C}$, from every 2D mask $k$ and compute for the frequency $\mathbf{r}^{\text{3D}}\in\mathbb{N}^{N}$ of every view as follows:
-$\mathbf{F}^{\text{3D}}=\sum_{t,k}\lambda^{k}_{t}*\mathbf{F}^{k,\text{2D}}_{t},\quad\quad\quad\quad\quad\quad\mathbf{r}^{\text{3D}}=\sum_{t,k}\lambda^{k}_{t},$ (4)
-where * is the outer product. The normalized pointwise 3D visual tokens are then obtained as follows:
+We set the visibility value $\lambda_{t,i}^{k}$ of any points whose pixel projections fall outside the $k$-th 2D mask’s boundaries or the disparity between projected depth $d$ and the collected depth $\mathbf{D}$ exceeds a defined depth threshold $\tau_{\text{depth}}$, or $|d_{i,t}-\mathbf{D}_{t}[\lfloor u_{i}\rfloor,\lfloor v_{i}\rfloor]|>\tau_{depth}$, to 0.
+Then, we accumulate 3D visual tokens $\mathbf{F}^{\text{3D}}\in\mathbb{R}^{N\times E\times C}$, from every 2D mask $k$ and compute for the frequency $\mathbf{r}^{\text{3D}}\in\mathbb{N}^{N}$ of every view as follows:
+$\mathbf{F}^{\text{3D}}=\sum_{t,k}\lambda^{k}_{t}*\mathbf{F}^{k,\text{2D}}_{t},\quad\quad\mathbf{r}^{\text{3D}}=\sum_{t,k}\lambda^{k}_{t},$ (2)
+where * is the outer product operation. The normalized pointwise 3D visual tokens are then obtained as follows:
 $\bar{\mathbf{F}}^{\text{3D}}_{i}=\begin{cases}\mathbf{F}^{\text{3D}}_{i}/\mathbf{r}^{\text{3D}}_{i}&\text{if }\mathbf{r}^{\text{3D}}_{i}>0\\ \mathbf{0}&\text{otherwise}\end{cases}.$ (3)
@@
 4 Experimental Results
-Datasets: We conducted experiments to assess the performance of the baselines and our proposed method using two common 3DIS datasets: ScanNet200 rozenberszki2022language and ScanNet++ yeshwanthliu2023scannetpp .
-The ScanNet200 dataset builds on the original ScanNet dai2017scannet by expanding its semantic categories from 20 to 200.
+Datasets: We conducted experiments to assess the performance of the baselines and our proposed method using two common 3DIS datasets: ScanNet200 [49] and ScanNet++ [56].
+The ScanNet200 dataset builds on the original ScanNet [17] by expanding its semantic categories from 20 to 200.
 Its instance segmentation benchmark includes 1,201 training scenes and 312 validation scenes with 198 object categories, significantly enriching the vocabulary and enhancing its capability for detailed 3D semantic and instance segmentation.
@@
 […]
-Implementation details: Following Open3DIS nguyen2023open3dis , we generate class-agnostic 3D proposals from ISBNet ngo2023isbnet pretrained on ScanNet200 or by lifting 2D masks to 3D using Detic detic .
-For RAM++ zhang2023recognize , we employ the Swin-L model, trained on a 14-million image dataset with an image size of 384px and a tagging threshold of 0.68.
-For LLM, we leverage Vicuna-7B vicuna2023 , fine-tuned for open-ended 2D Instance Segmentation yu2023towards .
-Table 1: Comparative results on the ScanNet200 dataset. Shaded text indicates a reference method, not a direct comparison. ‘-’ indicates results are not provided. The best results are in bold.
-[…]
-Mask3D Schult23mask3d 26.9 36.2 41.4 39.8 21.7 17.9 -
-Fully-sup 3DIS ISBNet ngo2023isbnet 24.5 32.7 37.6 38.6 20.5 12.5 -
-OpenMask3D takmaz2023openmask3d 15.4 19.9 23.1 17.1 14.1 14.9 -
-OV-3DIS Open3DIS nguyen2023open3dis 23.7 29.4 32.8 27.8 21.2 21.8 -
-Large-Vocab 21K DETIC classes (Ours) 8.5 11.7 13.1 9.9 7.2 8.3 52.2
-Image-Tagging RAM++ zhang2023recognize (Ours) 10.7 14.3 16.0 11.6 11.0 9.3 60.5
-Maskwise OSM yu2023towards (Ours) 14.4 19.8 23.9 18.9 13.5 10.2 65.0
-Pointwise Ours 16.0 22.0 24.7 20.0 14.3 13.2 65.2
-Table 2: Comparative results on the ScanNet++ dataset. Shaded text indicates a reference method, not a direct comparison. ‘-’ indicates results are not provided. The best results are in bold.
-Setting Method AP AP50 AP25 AR RC50 RC25 OE
-PointGroup jiang2020pointgroup 8.9 14.6 38.9 - - - -
-Fully-sup 3DIS SoftGroup vu2022softgroup 16.7 29.7 21.0 - - - -
-OpenMask3D takmaz2023openmask3d 2.0 2.7 3.4 4.6 8.3 12.4 -
-OV-3DIS Open3DIS (known classes) nguyen2023open3dis 13.1 20.8 24.6 22.1 33.9 39.1 -
-Large-Vocab 21K DETIC classes (Ours) 7.3 11.9 15.2 13.3 20.3 23.6 42.0
-Image-Tagging RAM++ zhang2023recognize (Ours) 9.1 15.5 19.1 16.0 24.8 28.7 48.5
-Maskwise OSM yu2023towards (Ours) 16.3 24.8 29.0 22.2 32.0 36.0 51.2
-Pointwise Ours 18.4 29.4 33.6 23.3 35.2 39.3 53.9
+Implementation details: Following Open3DIS [39], we generate class-agnostic 3D proposals from ISBNet [37] pretrained on ScanNet200 or by lifting 2D masks to 3D using Detic [63].
+For RAM++ [29], we employ the Swin-L model, trained on a 14-million image dataset with an image size of 384px and a tagging threshold of 0.68.
+For LLM, we leverage Vicuna-7B [9], fine-tuned for open-ended 2D Instance Segmentation [59].
+Setting Method ScanNet200 ScanNet++
+AP AP50 AP25 $_{\text{h}} AP$$_{\text{c}}$ $_{\text{t}} AP$ AP50 AP25 AR RC50 RC25
+Mask3D [50] 26.9 36.2 41.4 39.8 21.7 17.9 8.9 14.6 38.9 - - -
+Fully-sup 3DIS ISBNet [37] 24.5 32.7 37.6 38.6 20.5 12.5 16.7 29.7 21.0 - - -
+OpenMask3D [51] 15.4 19.9 23.1 17.1 14.1 14.9 2.0 2.7 3.4 4.6 8.3 12.4
+OV-3DIS Open3DIS [39] 23.7 29.4 32.8 27.8 21.2 21.8 13.1 20.8 24.6 22.1 33.9 39.1
+Large-Vocab 21K DETIC classes (Ours) 8.5 11.7 13.1 9.9 7.2 8.3 7.3 11.9 15.2 13.3 20.3 23.6
+Image-Tagging RAM++ [29] (Ours) 10.7 14.3 16.0 11.6 11.0 9.3 9.1 15.5 19.1 16.0 24.8 28.7
+Maskwise OSM [59] (Ours) 14.4 19.8 23.9 18.9 13.5 10.2 16.3 24.8 29.0 22.2 32.0 36.0
+Pointwise Ours 16.0 22.0 24.7 20.0 14.3 13.2 18.4 29.4 33.6 23.3 35.2 39.3
+Table 1: Comparative results on the ScanNet200 and ScanNet++ datasets.
+Shaded text indicates a reference method, not a direct comparison.
+‘-’ indicates results are not provided.
+The best results are in bold.
 4.1 Comparison with Baselines
-We compare our approach to the proposed baselines using the OE-3DIS setting on the ScanNet200 (Tab.
-1) and ScanNet++ (Tab.
-2) datasets.
-For reference, we also present results from OV-3DIS methods, including OpenMask3D takmaz2023openmask3d and Open3DIS nguyen2023open3dis , as well as fully-supervised methods like ISBNet ngo2023isbnet and Mask3D Schult23mask3d for ScanNet200; and PointGroup jiang2020pointgroup and SoftGroup vu2022softgroup for ScanNet++.
-For Scannet200: We obtain class-agnostic 3D proposals from two sources: 3D masks from a 3DIS network such as ISBNet ngo2023isbnet , and 2D Lift 3D masks from Open3DIS nguyen2023open3dis .
-Our proposed approach outperforms other baselines in both AP and OE scores.
+We compare our approach to the proposed baselines using the OE-3DIS setting on the ScanNet200 and ScanNet++ datasets in Tab.
+1.
+For reference, we also present results from OV-3DIS methods, including OpenMask3D [51] and Open3DIS [39], as well as fully-supervised methods like ISBNet [37] and Mask3D [50] for ScanNet200; and PointGroup [30] and SoftGroup [52] for ScanNet++.
+For Scannet200: We obtain class-agnostic 3D proposals from two sources: 3D masks from a 3DIS network such as ISBNet [37], and 2D Lift 3D masks from Open3DIS [39].
+Our proposed approach outperforms other baselines in the AP score.
 Furthermore, the performance progression of the baselines, in the specified order, clearly justifies the motivation behind each baseline compared to its predecessor, as discussed in Sec.
 3.3.
-Interestingly, our approach also surpasses OpenMask3D takmaz2023openmask3d (16.0 vs.
+Interestingly, our approach also surpasses OpenMask3D [51] (16.0 vs.
 15.4 in AP), even though OpenMask3D utilizes provided class names.
@@
 For ScanNet++: Due to the extensive scale and vast array of classes in ScanNet++, the performance of 3D mask results from ISBNet is inadequate.
-Consequently, we solely rely on the utilization of 2D Lift 3D masks from Open3DIS nguyen2023open3dis .
+Consequently, we solely rely on the utilization of 2D Lift 3D masks from Open3DIS [39].
 We notice a consistent trend akin to the results observed in ScanNet200.
-Particularly noteworthy is the significant outperformance of our approach compared to OV-3DIS methods or even fully-supervised 3DIS, as evidenced by higher AP scores.
-This underscores the promising application of OE-3DIS in navigating complex 3D scene structures.
+Particularly noteworthy is the significant outperformance of our approach compared to OV-3DIS methods or even fully supervised 3DIS, as evidenced by higher AP scores.
+This underscores the promising application of OE-3DIS in navigating complex 3D scene.
 Qualitative comparison: In Fig.
-3, both our proposed baselines and approach effectively assign class names to 3D proposals in OE-3DIS settings.
-The first column of ScanNet200 qualitative results shows our architectures accurately predicting all 3D proposals.
-The second column highlights differences between our approach and the baselines.
-Our method precisely identifies the ‘painting’ object as “wall painting”, while other baselines predict less accurate object labels.
-In the final column, our 3D mask proposals offer more detail than ground truth, enabling predictions of classes not in ScanNet++’s vocabulary, such as “photocopier”.
-Figure 3: Qualitative examples are provided for ScanNet200 dai2017scannet (first two columns) and ScanNet++ yeshwanthliu2023scannetpp (last two columns).
-Both our baselines and approach yield notably good results, particularly in terms of accurately identifying class names even though they do not match exactly the GT classes.
+4, both our proposed approach and baseline methods effectively assign class names to 3D proposals under OE-3DIS settings.
+The first column of qualitative results on ScanNet200 demonstrates that our architectures correctly predict class labels for all 3D proposals.
+The second column highlights notable differences, with our method accurately identifying the ‘painting’ object as a ‘wall painting’, whereas the baselines produce less precise labels.
+In the final column, our detailed 3D mask proposals surpass the granularity of the ground-truth annotations, enabling accurate recognition of classes absent from ScanNet++’s vocabulary, such as ‘photocopier’.
+This illustrates how OE-3DIS allows the model to comprehensively understand a 3D scene without restricting to fixed categories.
+[…]
+L2 Norm (Open3DIS) 5.7 8.2 9.8 6.8 4.9 5.2
+Memory Fusion (OVIR-3D) 8.4 11.6 14.4 7.2 7.0 11.6
+Max 6.3 8.4 10.2 8.5 4.7 5.6
+Random 13.2 18.9 21.6 17.7 12.3 10.0
+Mean 14.5 20.1 22.6 18.2 13.1 11.2
+Weighted Average (Ours) 16.0 22.0 24.7 20.0 14.3 13.2
 Table 2: Ablation on point aggregation techniques
 […]
-Max 6.3 8.4 10.2 8.5 4.7 5.6 56.3
-Random 13.2 18.9 21.6 17.7 12.3 10.0 60.4
-Mean 14.5 20.1 22.6 18.2 13.1 11.2 63.1
-Weighted Average 16.0 22.0 24.7 20.0 14.3 13.2 65.2
+3D masks DETIC 11.7 16.0 16.3 10.3 8.1
+2D Lift 3D masks DETIC 11.7 18.6 11.3 12.1 11.8
+3D + 2D Lift 3D masks DETIC 16.0 22.0 20.0 14.3 13.2
+3D + 2D Lift 3D masks SAM 15.4 20.5 19.2 14.8 11.7
 Table 3: Study on different types 3D proposals.
 […]
-3D masks DETIC 11.7 16.0 18.3 16.3 10.3 8.1 56.4
-2D Lift 3D masks DETIC 11.7 18.6 23.5 11.3 12.1 11.8 54.3
-3D + 2D Lift 3D masks DETIC 16.0 22.0 24.7 20.0 14.3 13.2 65.2
-3D + 2D Lift 3D masks SAM 15.4 20.5 22.9 19.2 14.8 11.7 64.8
+BERT [20] 13.5 19.0 21.2 17.6 11.6 11.0
+CLIP [43] 15.9 22.0 24.5 20.3 13.7 13.3
+Sentence Transformer [44] 16.0 22.0 24.7 20.0 14.3 13.2
+Table 4: Study on different text encoders in evaluation metrics
+Modules # Params(M) FLOPs (G) Time (s)
+Pw. CLIP extract (CLIP) 427.94 191.11 97
+QFormer 185.66 2.74 250
+LLM (Vicuna 7B) 6,738.42 850.01 50
+RAM++ 329.49 104.43 40
+Generate 3D proposals N/A 357.99 152
+Pw. Visual Token Lift (PVTL) N/A 4.915 314
+Table 5: Expected latency, FLOPs, and runtime of each module for a 3D scene. Pw. denotes ‘Pointwise’. Generating 3D proposals from 2D and Pw. visual token lifting are non-parametric modules
+Figure 4: Qualitative examples are provided for ScanNet200 [17] (first two columns) and ScanNet++ [56] (last two columns).
+Both our baselines and approach yield notably good results, particularly in terms of accurately identifying class names even though they do not match exactly the GT classes.
+Method AP Total Time (s) RAM++ Generate 3D proposals CLIP QFormer PVTL LLM
+Open3DIS N/A 249
+Large Vocab 8.5 249
+Image Tagging 10.7 289
+Maskwise 14.4 452
+Pointwise 16.0 766
+Table 6: Expected runtime for our proposed baselines and methods with the Open3DIS for a sample 3D scene
+Text Prompt AP
+“What is in the segmentation mask? Assistant:” 16.0
+“Can you describe what is in the segmentation mask region? Assistant:” 15.3
+“What can you see in the segmentation mask region? Assistant:” 15.9
+“Could you use a few words to describe what is in the segmentation mask? Assistant:” 15.2
+“What is this segmentation mask? Assistant:” 15.8
+Table 7: Study on different text prompts used to query the LLM.
 4.2 Ablation Study
-To study many design choices of our pointwise approach, we intensively carry out ablation study on the ScanNet200 dai2017scannet dataset.
+To study many design choices of our pointwise approach, we intensively carry out ablation study on the ScanNet200 [17] dataset.
 Study on different point feature aggregation techniques (Eq.
-6) is shown in Tab.
-3.
-We evaluate four techniques for combining point features from a given 3D mask: max, random (randomly selecting one point), mean, and weighted average (our proposed operation).
+4) is shown in Tab.
+2.
+We evaluate four techniques for combining point features from a given 3D mask: L2-norm, max, random (randomly selecting one point), mean, and weighted average (our proposed operation).
 The weighted average technique achieved the highest performance with an AP score of 16.0, outperforming the other methods.
@@
 As reported in Tab.
-4, combining 3D proposals from both 3D masks and 2D Lift 3D masks yielded the most favorable results.
+3, combining 3D proposals from both 3D masks and 2D Lift 3D masks yielded the most favorable results.
 This outcome validates our choice of 3D proposals for the ScanNet200 dataset.
 Additionally, the 2D Lift 3D proposals obtained from DETIC demonstrated slightly superior performance compared to those from SAM.
-This observation diverges from the methodology adopted in Open3DIS nguyen2023open3dis , where provided class names are utilized to employ Grounded-SAM ren2024grounded instead of SAM sam , as in our OE-3DIS scenario.
+This observation diverges from the methodology adopted in Open3DIS [39], where provided class names are utilized to employ Grounded-SAM [45] instead of SAM [31], as in our OE-3DIS scenario.
 Study on different text encoders for evaluation metrics is described in Tab.
-5.
-We assess three text decoders: BERT devlin2018bert , CLIP radford2021learning , and Sentence Transformer sentence-transformer .
-Among them, the Sentence Transformer embeddings attain the highest AP scores, whereas the CLIP text embeddings achieve the highest OE score.
-The OE score indicates that the text embeddings from CLIP exhibit a higher degree of similarity to each other, contrasting with those from the Sentence Transformer, which tend to describe complex sentences rather than solely focusing on a single class name.
-Hence, we lean towards selecting the Sentence Transformer as our recommended text encoder.
+4.We evaluate three text decoders: BERT [20], CLIP [43], and Sentence Transformer [44].
+Among these, Sentence Transformer embeddings achieve the highest AP scores, outperforming CLIP and BERT, which yield comparatively lower scores.
+We also observe that CLIP embeddings exhibit greater similarity among themselves compared to those from BERT.
+Moreover, Sentence Transformer embeddings are better suited for capturing complex sentence-level descriptions rather than focusing solely on individual class names.
+Therefore, we recommend Sentence Transformer as our preferred text encoder.
+Latency analysis of our baselines and proposed method is presented in Tab.
+5 (excluding the runtime for generating 2D proposals).
+Additionally, Tab.
+6 compares the total runtime across our baselines, proposed methods, and the state-of-the-art 3DIS approach, Open3DIS.
+As shown, our method prioritizes accuracy at the cost of increased runtime, meaning higher accuracy comes with slower execution.
 Study on different text prompts used to query the LLM: We experiment with various input text prompts to query the LLM for class names.
@@
 Assistant:” from the variants for our approach.
-Table 5: Study on different text encoders used in evaluation metrics
-[…]
-BERT devlin2018bert 13.5 19.0 21.2 17.6 11.6 11.0 70.6
-CLIP radford2021learning 15.9 22.0 24.5 20.3 13.7 13.3 71.6
-Sentence Transformer sentence-transformer 16.0 22.0 24.7 20.0 14.3 13.2 65.2
-Table 6: Study on different text prompts used to query the LLM.
-Text Prompt AP OE
-“What is in the segmentation mask? Assistant:” 16.0 65.2
-“Can you describe what is in the segmentation mask region? Assistant:” 15.3 64.3
-“What can you see in the segmentation mask region? Assistant:” 15.9 65.0
-“Could you use a few words to describe what is in the segmentation mask region? Assistant:” 15.2 64.0
-“What is this segmentation mask? Assistant:” 15.8 65.0
 5 Discussion and Conclusion
-Limitations: Though our method enhances open-ended 3D scene understanding, it has notable drawbacks.
-Firstly, it heavily relies on 2D visual tokens from the pretrained OSM, which are trained on instance segmentation datasets with limited classes.
-This hampers the model’s ability to recognize a wide range of classes in an open-world environment.
-Secondly, the performance of OE-3DIS is contingent upon class-agnostic 3D instance segmentation (3DIS) methods, which in turn rely on the quality of 3D representation and 2D-3D mapping, such as camera calibrations and depth images.
+Limitations: The motivation behind our proposed Open-ended 3D Instance Segmentation (OE-3DIS) arises from limitations in current Open-Vocabulary 3D Point Cloud Instance Segmentation (OV-3DIS) methods, which still depend on a predefined set of class names during testing.
+This constraint is impractical in scenarios lacking prior knowledge of class names, such as a robot navigating unfamiliar environments.
+Although our method advances open-ended 3D scene understanding, it still faces notable limitations.
+Firstly, our approach heavily depends on 2D visual tokens extracted from a pretrained OSM, which itself is trained on instance segmentation datasets containing only a limited number of classes.
+Consequently, this restricts the model’s capability to recognize an extensive range of classes in truly open-world contexts.
+Secondly, the performance of OE-3DIS relies significantly on class-agnostic 3DIS methods, whose effectiveness in turn depends on the accuracy of 3D representations and the quality of 2D-to-3D mapping, including factors like camera calibration and depth image quality.
 Conclusion: We have introduced Open-Ended 3D Point Cloud Instance Segmentation (OE-3DIS), which generates 3D masks and object class names without predefined labels during testing.
 We have explored baselines using OV-3DIS methods and MLLMs, and introduced a pointwise training-free approach leveraging OSM.
-We have also proposed the OE score, tailored for OE-3DIS.
 Experiments on ScanNet200 and ScanNet++ show our approach’s superior performance, notably outperforming Open3DIS (SOTA on OV-3DIS) on ScanNet++ without ground-truth class names.